# Aligning Speech to Languages to Enhance Code-switching Speech Recognition

### v2

@@ V-D Comparison with state-of-the-art methods
We compared the performance of the proposed method against state-of-the-art approaches, including train-from-scratch and pre-trained models. These methods are evaluated on both model parameters and the CS-ASR performance is outlined in Tables VII. For the ASRU dataset, the hybrid CTC/attention model incorporating LAL with language weights achieves the best MER of 11.7% among the train-from-scratch approaches under consideration. While the bi-encoder and LPB methods exhibit higher performance compared to the baseline model, these approaches require an additional Conformer encoder and a Transformer decoder, respectively, resulting in a notable increase in model parameters.
@@
The proposed method has also been validated through fine-tuning the Whisper model. The Whisper model initially demonstrates low zero-shot performance on both datasets, with particularly high error rates of 90.8% and 69.7% observed on the SEAME test sets. However, models fine-tuned on the respective training sets demonstrate improved performance over the pre-trained Whisper, outperforming models trained from scratch. Fine-tuning with the proposed LAL leads to a higher performance compared to pure fine-tuning. This is consistent with the results observed in the hybrid CTC/attention model, highlighting the effectiveness of the proposed method.
Fine-tuning Whisper with either LID or LAL consistently results in improved performance compared to standard fine-tuning, indicating that incorporating auxiliary frame-level LID enhances CS-ASR performance. Furthermore, the Whisper model fine-tuned with the proposed LAL outperforms models that rely on auxiliary LID using language labels derived from FA-LB. This observation supports our statement in Section III-B that the proposed LAL integrates the CS-ASR and LID into a mutual optimization process during training, leading to better performance than conventional LID approaches. However, the proposed LAL shows lower performance than jointly optimizing ASR and LID with FA-UB. This implies that the proposed LAL generates better pseudo labels compared to force alignment computed by models with comparable performance levels, though these labels may still fall short of the accuracy provided by highly precise language labels, such as force alignment computed by fine-tuned Whisper-large model and gold-standard annotations.
@@
Ling. hint type $\text{test}_{\texttt{man}}$ $\text{test}_{\texttt{sge}}$ ASRU test
MER $\downarrow$ Acc. $\uparrow$ MER $\downarrow$ Acc. $\uparrow$ MER $\downarrow$ Acc. $\uparrow$
@@
TABLE IX: Performance evaluation of prompting LLM with generative error correction after incorporating the linguistic hint by employing MER (%). “$ft$” denotes fine-tuning, “$gt$” denotes the linguistic hint using ground-truth utterance-level language label, and “$pred$” denotes the predicted linguistic hint
@@ V-E Incorporating LLM with linguistic hint
Similar to existing works [50, 8], it is not surprising that the LID accuracy achieved with the incorporation of textual information is higher than the by-product of the proposed LAL. This can be attributed to the fact that an ASR system models both acoustic and language characteristics, whereas an LID system generally focuses solely on acoustic information. Furthermore, a decrease in LID accuracy often results in a corresponding degradation in CS-ASR performance of the generated linguistic hints. The prompt incorporating the linguistic hint, which serves as the ground-truth language label, therefore exhibits significantly higher performance on the SEAME and ASRU datasets compared to other prompts.
@@
The performance of CS-ASR systems with external language modeling are summarized in Table IX. We observe that incorporating an LM via shallow fusion improves the CS-ASR performance. The use of LLM does not lead to performance improvement on the SEAME dataset unless accompanied by the ground-truth linguistic hint. In contrast, the LLM can improve the performance on the ASRU dataset, achieving an MER of 11.0% for all types of prompts. We analyzed the performance by considering the data distribution as shown in Table II. Since all utterances in the ASRU test set are code-switched, these predicted linguistic hints all exhibit high LID accuracy, and thus show comparable CS-ASR performance.
@@
Ling. hint type $\text{test}_{\texttt{man}}$ $w/o$ $interjections$
@@
no hint 16.4 - 13.7 -
LID output (LAL) 17.0 80.3 14.5 81.1
1st hypo. 16.5 93.3 13.6 95.4
hypos. vote 16.8 93.2 13.7 95.3
LAL & hypos. vote 16.8 92.9 13.7 95.3
groundtruth 15.7 - 13.4 -
@@ V-F Impact of interjections
As described in Section IV-A, the SEAME dataset was collected from Singapore and Malaysia, where code-switching is more frequent than the ASRU dataset—interjections such as “lah”, “lor”, and “ya” that often occur in the former dataset. Annotating interjections is a challenging task even for individuals with bilingual expertise. Hence, interjections can introduce confusion during language modeling, leading to a degradation in the CS-ASR performance for LLMs that have not been trained on them. We therefore removed interjections from utterances in the SEAME training and $\text{test}_{\texttt{man}}$ without changing their semantic information. The CS-ASR performance and the LID accuracy of the linguistic hint prediction before and after removing interjections when incorporating LLM are shown in Table X. In addition, the first row shows the MER of the hybrid CTC/attention model with LAL. This is computed by comparing the decoded hypotheses and ground-truth transcriptions after removing interjections within them.
@@ V-G Language identification performance
Our primary target is to improve the CS-ASR performance of the model with the assistance of language information instead of achieving high LID performance. Although the proposed LAL generally improves the CS-ASR performance, tuning an appropriate $\beta$ for the LAL loss term in (15) is crucial for optimal performance improvement. Therefore, we present the LID performance of models trained with various $\beta$ values in Table XI.
As shown in the table, the model achieves the lowest LID performance when $\beta=5.0$. This suggests that using a higher $\beta$ value may not necessarily achieve the optimal LID performance. The highest LID performance is achieved with $\beta=1.5$. This indicates that the model optimized with an appropriate $\beta$ value can achieve high performance for both CS-ASR and LID tasks. In addition, this implies that higher ASR performance can benefit the LID branch.
@@ VI-A Frame-to-token alignment
CTC. In the hybrid CTC/attention ASR model, the frame-to-token alignment can be computed from both CTC and the cross-attention process within the ASR decoder. However, it is worth noting that CTC predictions contain $<$blank$>$ token that results in peaky behaviour [52, 53]. Since the $<$blank$>$ token lacks a language attribute and is unable to be converted into a language label, this peaky behavior leads to fewer language labels in the pseudo label sequences compared to those derived through the ASR decoder. Therefore, we use the frame-to-token alignment computed from the cross-attention attention weight matrix within the ASR decoder but not the CTC to generate pseudo language labels.
Force alignment. One can also use additional processing such as force alignment to retrieve token-level timestamps and achieve frame-to-token alignment. However, state-of-the-art multilingual ASR models, such as Whisper, require fine-tuning to achieve satisfactory performance. As presented in Table VII, the pre-trained Whisper-small model achieves MERs of 24.9%, 90.8%, and 69.7% on the ASRU and SEAME test sets, respectively. As a result, applying force alignment to code-switching speech using pre-trained multilingual ASR models is undesirable.
Results in Table VII show that the proposed LAL method outperforms performing LID with force-alignment language labels computed by the fine-tuned Whisper-small model (FA-LB). Besides, force alignment incurs a substantial computational burden, as it must be applied to the entire training dataset using fine-tuned Whisper models. While LID with FA-UB exhibits higher performance than the proposed LAL, it results in an even greater computational cost than FA-LB due to reliance on the Whisper-large model. This contradicts the motivation of our work, which is to enrich the CS-ASR model with language information without the need for an additional annotating process.
TABLE XII: Comparison of the original ASR with LAL outputs, the LM late fusion outputs, the LLM-generated outputs, and the ground-truth transcripts for the SEAME dataset, errors and corrected words are marked in red and blue, respectively.
@@
+ LLM LoRA $ft$ + ling. hint $(gt)$ ah yah
@@ VI-B Where the errors happen
The errors have been analyzed to gain insights into factors affecting the ASR performance. Compared to deletion and insertion rates, a significantly higher substitution rate is observed from the aforementioned results. The high substitution rate impose a challenge in language modeling due to a larger vocabulary and language confusion arising from a code-switching scenario. Given that code-switched text may be more readily achieved than speech, developing a robust code-switching language model is desirable to address language confusion effectively.
@@
To gain further insights into the CS-ASR performance on the SEAME dataset, we compared the decoded outputs of the ASR model incorporating LAL, and its counterparts enhanced by LM and LLM respectively. The results are juxtaposed with the ground-truth transcription in Table XII. Here, errors are marked in red and the tokens corrected by LM and LLM are marked in blue. One factor that contributes to the errors in the CS-ASR task is language confusion. For instance, the expression “ah yeah” shares both the pronunciation and semantic characteristics with its Chinese equivalent “唉 呀”. Another notable factor that contributes to errors is the liaison, where two Mandarin characters or English words can erroneously be classified as a single entity. The liaison effect can be particularly pronounced in a spontaneous code-switching speech signal.
The LLM fine-tuned on the SEAME training set demonstrated effectiveness in correcting grammatical errors. Words such as “happened” are consequently adjusted to the correct tense. However, the model is less adept at accommodating colloquial expressions. Therefore, the third and fourth cases illustrated in Table XII have been modified to a more formal expression, leading to a higher mixed error rate.
@@ VI-C Attention matrices
We visualize the multi-head attention matrices within various ASR decoder layers corresponding to a speech sample as shown in Fig. 6, where the x- and y-axis denote speech frames (encoder output) and tokens, respectively. The model is the best-performing CS-ASR model with the proposed LAL on the ASRU data.
Existing works have shown that shallow layers within a Transformer-based model focus on low-level speech information such as phonetics and language, while deep layers often capture high-level information such as semantic and sequential characteristics [54, 55]. This is consistent with the visualization, motivating us to employ the last decoder layer. Therefore, the average of cross-attention matrices within the last decoder is used to compute the frame-level pseudo-language labels.
Specifically, compared to the attention matrices in the third decoder layer, those in the last decoder layer capture sequential information more effectively and display a monotonic mapping of frames and tokens. However, Fig. 6(b) shows that the monotonic mapping can be blurred due to the difficulty in learning the attention pattern of each attention head during training. To mitigate this challenge, the averaged frame-to-token alignment across all attention heads is used instead of relying on a single attention head.
@@ VII Conclusion
We proposed to align speech to languages to enhance the CS-ASR performance in both ASR and LLM-based processing. For the ASR model, we introduced a language alignment loss (LAL) to enrich the model with language information. Models equipped with the proposed LAL consistently exhibited higher CS-ASR performance than the vanilla configuration, with only a negligible increase in model parameters during training. In addition, the proposed LAL outperforms frame-level LID using language labels obtained through force alignment. These demonstrated the effectiveness of the proposed LAL. After incorporating language weights in LAL to address language imbalance in the ASRU data during training, the proposed method obtained further performance improvement. Beyond ASR, we proposed leveraging a linguistic hint, which is derived from LAL outputs and decoded hypotheses, to guide the prompting and enhance the LLM-based generative error correction. Experimental results indicate that an accurate linguistic hint can significantly improve the CS-ASR performance in scenarios involving both monolingual and code-switching utterances. Finally, the errors within the hypotheses are analyzed. The LLM fine-tuned on the SEAME data has shown to be effective in correcting grammatical errors, which, in contrast, leads to a lower ASR performance for spontaneous and colloquial speech.

### v3

@@ V-D Comparison with state-of-the-art methods
We compare the performance of the proposed method against state-of-the-art approaches, including train-from-scratch and pre-trained models. These methods are evaluated on both model parameters and the CS-ASR performance is outlined in Table VII. For the ASRU dataset, the hybrid CTC/attention model incorporating LAL with language weights achieves the best MER of 11.7% among the train-from-scratch approaches under consideration. While the bi-encoder and LPB methods exhibit higher performance compared to the baseline model, these approaches require an additional Conformer encoder and a Transformer decoder, respectively, resulting in a notable increase in model parameters.
@@
The proposed method has also been validated through fine-tuning the Whisper model. The Whisper model initially suffers from low zero-shot performance on both datasets, with particularly high error rates of 90.8% and 69.7% observed on the SEAME test sets. However, models fine-tuned on the respective training sets demonstrate improved performance over the pre-trained Whisper, outperforming models trained from scratch. Fine-tuning with the proposed LAL leads to a higher performance compared to pure fine-tuning. This is consistent with the results observed in the hybrid CTC/attention model, highlighting the effectiveness of the proposed method.
Fine-tuning Whisper with either LID or LAL consistently results in improved performance compared to standard fine-tuning, indicating that incorporating auxiliary frame-level LID enhances CS-ASR performance. Furthermore, the Whisper model fine-tuned with the proposed LAL outperforms models that rely on auxiliary LID using language labels derived from FA-LB. This observation supports our statement in Section III-B that the proposed LAL integrates the CS-ASR and LID into a mutual optimization process during training, leading to better performance than conventional LID approaches. However, the proposed LAL shows lower performance than jointly optimizing ASR and LID with FA-UB. This implies that the proposed LAL generates better pseudo labels compared to forced alignment computed by models with comparable performance levels, though these labels may still fall short of the accuracy provided by highly precise language labels, such as forced alignment computed by the fine-tuned Whisper-large model and gold-standard annotations.
@@
LLM LoRA ftft $\text{test}_{\texttt{man}}$ $\text{test}_{\texttt{sge}}$ ASRU test
Ling. hint type MER $\downarrow$ Acc. $\uparrow$ MER $\downarrow$ Acc. $\uparrow$ MER $\downarrow$ Acc. $\uparrow$
@@
TABLE IX: Performance evaluation of prompting LLM with generative error correction after incorporating the linguistic hint by employing MER (%). “ftft” denotes fine-tuning, “gtgt” denotes the linguistic hint using ground-truth utterance-level language label, and “predpred” denotes the predicted linguistic hint
@@ V-E Incorporating LLM with linguistic hint
Similar to [52, 10], it is not surprising that the LID accuracy achieved with the incorporation of textual information is higher than the by-product of the proposed LAL. This can be attributed to the fact that an ASR system models both acoustic and language characteristics, whereas an LID system generally focuses solely on acoustic information. Furthermore, a decrease in LID accuracy often results in a corresponding degradation in CS-ASR performance of the generated linguistic hints. The prompt incorporating the linguistic hint, which serves as the ground-truth language label, therefore exhibits significantly higher performance on the SEAME and ASRU datasets compared to other prompts.
@@
The performance of CS-ASR systems with external language modeling is summarized in Table IX. We observe that incorporating an LM via shallow fusion improves the CS-ASR performance. The use of LLM does not lead to performance improvement on the SEAME dataset unless accompanied by the ground-truth linguistic hint. In contrast, LLM can improve the performance on the ASRU dataset, achieving an MER of 11.0% for all types of prompts. We analyzed the performance by considering the data distribution as shown in Table II. Since all utterances in the ASRU test set are code-switched, these predicted linguistic hints all exhibit high LID accuracy, and thus show comparable CS-ASR performance.
@@
Method $\text{test}_{\texttt{man}}$ $w/o$ interjectionsinterjections
@@
+ LLM LoRA $ft$ (no hint) 16.4 - 13.7 -
w/ LID output (LAL) 17.0 80.3 14.5 81.1
w/ 1st hypo. 16.5 93.3 13.6 95.4
w/ hypos. vote 16.8 93.2 13.7 95.3
w/ LAL & hypos. vote 16.8 92.9 13.7 95.3
w/ groundtruth 15.7 - 13.4 -
@@ V-F Impact of interjections
As described in Section IV-A, the SEAME dataset was collected from Singapore and Malaysia, where code-switching is more frequent than the ASRU dataset—interjections such as “lah”, “lor”, and “ya” often occur in the former dataset. Annotating interjections is a challenging task even for individuals with bilingual expertise. Hence, interjections can introduce confusion during language modeling, leading to a degradation in the CS-ASR performance for LLMs that have not been trained on them. We therefore removed interjections from utterances in the SEAME training and $\text{test}_{\texttt{man}}$ without changing their semantic information. The CS-ASR performance and the LID accuracy of the linguistic hint prediction before and after removing interjections when incorporating LLM are shown in Table X. In addition, the first row shows the MER of the hybrid CTC/attention model with LAL. This is computed by comparing the decoded hypotheses and ground-truth transcriptions after removing interjections within them.
@@ V-G Language identification performance
Our primary objective is to improve the CS-ASR performance of the model with the assistance of language information instead of achieving high LID performance. Although the proposed LAL generally improves the CS-ASR performance, tuning an appropriate $\beta$ for the LAL loss term in (15) is crucial for optimal performance improvement. Therefore, we present the LID performance of models trained with various $\beta$ values in Table XI.
As shown in the table, the model achieves the lowest LID performance when $\beta=5.0$. This suggests that using a higher $\beta$ value may not necessarily achieve the optimal LID performance. The highest LID performance is achieved with $\beta=1.5$. This indicates that the model optimized with an appropriate $\beta$ value can achieve high performance for both CS-ASR and LID tasks. In addition, this implies that higher ASR performance can improve the LID branch.
@@ VI-A Frame-to-token alignment
CTC. In the hybrid CTC/attention ASR model, the frame-to-token alignment can be computed from both CTC and the cross-attention process within the ASR decoder. However, CTC predictions contain <<blank>> token that results in peaky behavior [54, 55]. Since the <<blank>> token lacks a language attribute and cannot be converted into a language label, this peaky behavior leads to fewer language labels in the pseudo label sequences compared to those derived through the ASR decoder. Therefore, we use the frame-to-token alignment computed from the cross-attention weight matrix within the ASR decoder but not from CTC to generate pseudo-language labels.
Forced alignment. We can also use additional processing (e.g., forced alignment) to retrieve token-level timestamps and achieve frame-to-token alignment. However, state-of-the-art multilingual ASR models, such as Whisper, require fine-tuning to achieve satisfactory performance. As presented in Table VII, the pre-trained Whisper-small model achieves MERs of 24.9%, 90.8%, and 69.7% on the ASRU and SEAME test sets, respectively. As a result, applying forced alignment to code-switching speech using pre-trained multilingual ASR models is undesirable.
Results presented in Table VII show that the proposed LAL method outperforms applying LID with forced-alignment language labels computed by the fine-tuned Whisper-small model (FA-LB). Moreover, forced alignment incurs a substantial computational burden, as it must be applied to the entire training dataset using fine-tuned Whisper models. While LID with FA-UB exhibits higher performance than the proposed LAL, it results in an even greater computational cost than FA-LB due to its reliance on the Whisper-large model. This contradicts the motivation of our work, which is to enrich the CS-ASR model with language information without the need for an additional annotating process.
TABLE XII: Comparison of the original ASR with LAL outputs, the LM late fusion outputs, the LLM-generated outputs with ground-truth linguistic hint, and the ground-truth transcripts. The first four examples are from the SEAME dataset, and the last one is from the ASRU dataset
@@
在 雨 田 机 场 体 验 了 一 下 shower room
+ LLM LoRA ftft ah yah
@@ VI-B Where the errors happen
The errors have been analyzed to gain insight into the factors affecting ASR performance. Compared to the deletion and insertion rates, a significantly higher substitution rate is observed from the aforementioned results. The high substitution rate imposes a challenge in language modeling due to a larger vocabulary and language confusion arising from a code-switching scenario. Given that code-switched text may be more readily achieved than speech, developing a robust code-switching language model is desirable to address language confusion effectively.
@@
To gain further insights into the CS-ASR performance, we compared the decoded outputs of the ASR model incorporating LAL and its counterparts enhanced by LM and LLM. The results are juxtaposed with the ground-truth transcription in Table XII. One factor that contributes to errors in the CS-ASR task is language confusion. For instance, the expression “ah yeah” shares both the pronunciation and semantic characteristics with its Chinese equivalent “唉 呀”. Another notable factor that contributes to errors is the liaison, where two Mandarin characters or English words can erroneously be classified as a single entity. The liaison effect can be particularly pronounced in a spontaneous code-switching speech signal.
The fine-tuned LLM demonstrated effectiveness in correcting grammatical errors for SEAME samples. Words such as “happened” are consequently adjusted to the correct tense. However, the model is less adept at accommodating colloquial expressions. Therefore, the third and fourth cases illustrated in Table XII have been modified to a more formal expression, leading to a higher mixed error rate. For the ASRU example, the fine-tuned LLM successfully corrects the incorrectly predicted entity name, likely due to its contextual understanding and world knowledge acquired during large-scale pre-training.
@@ VI-C Attention matrices
We visualize the multi-head attention matrices within various ASR decoder layers corresponding to a speech sample as shown in Fig. 6, where the abscissa and the ordinate axes denote speech frames (encoder output) and tokens, respectively. The model is the best-performing CS-ASR model with the proposed LAL on the ASRU data.
Existing works have shown that shallow layers within a Transformer-based model focus on low-level speech information such as phonetics and language, while deep layers often capture high-level information such as semantic and sequential characteristics [56, 57]. This is consistent with the visualization, motivating us to employ the last decoder layer. Therefore, the average of cross-attention matrices within the last decoder layer is used to compute the frame-level pseudo-language labels.
Specifically, compared to the attention matrices in the third decoder layer, those in the last decoder layer capture sequential information more effectively and display a monotonic mapping between frames and tokens. However, Fig. 6(b) shows that the monotonic mapping can be blurred due to the difficulty in learning the attention pattern of each attention head during training. To mitigate this challenge, the averaged frame-to-token alignment across all attention heads is used instead of relying on a single attention head.
VI-D LAL in Speech LLMs
The proposed method holds potential for integration into both the pre-training and fine-tuning stages of multi-modal LLMs, such as speech LLMs, that support speech input, particularly for code-switching scenarios.
The speech encoder in speech LLMs is typically derived from the encoder of a large-scale pre-trained ASR model [58, 59, 39]. Incorporating the proposed LAL into the speech encoder or tokenizer [39, 60] during pre-training on code-switching data may effectively enrich the model with code-switching capability. This integration can help the model better capture language boundaries and transitions, which are particularly prevalent in Southeast Asian multilingual contexts. Moreover, while large-scale ASR models already achieve strong recognition performance, their encoders produce rich speech representations that can support more accurate LID decisions for LAL. These LID decisions can then be integrated into the prompt, providing the LLM with explicit language guidance during text generation.
However, the proposed LAL may not be well-suited for general-purpose speech LLMs in which code-switching data rarely exists in the training corpus. In such settings, the supervision provided by LAL would largely resemble an auxiliary utterance-level language identification task, offering limited benefit. As a result, incorporating LAL into speech LLMs may not yield meaningful improvements in these scenarios.
Furthermore, the LLM module within speech LLMs can internally perform generative error correction before generating the final ASR output. This correction is achieved through the reasoning capabilities of LLMs, applied to either the N-best hypotheses or a single hypothesis during inference. Recent LLMs have demonstrated stronger multilingual capabilities in both understanding and reasoning compared to earlier models such as LLaMA-2 used in this work [61, 62]. These advancements can enhance internal error correction in code-switching ASR, particularly when guided by the proposed linguistic hint.
@@ VII Conclusion
We proposed to align speech to languages to enhance the CS-ASR performance in both ASR and LLM-based processing. For the ASR model, we introduced a language alignment loss (LAL) to enrich the model with language information. Models equipped with the proposed LAL consistently achieve higher CS-ASR performance than the vanilla configuration, with only a negligible increase in model parameters during training. In addition, the proposed LAL outperforms frame-level LID using language labels obtained through forced alignment. These demonstrated the effectiveness of the proposed LAL. After incorporating language weights into LAL to address language imbalance in the ASRU data during training, the proposed method obtained further performance improvement. Beyond ASR, we proposed leveraging a linguistic hint, which is derived from LAL outputs and decoded hypotheses, to guide the prompting and enhance the LLM-based generative error correction. Experimental results indicate that an accurate linguistic hint can significantly improve CS-ASR performance in scenarios involving both monolingual and code-switching utterances. Finally, the errors within the hypotheses are analyzed. The LLM fine-tuned on the SEAME data has shown effectiveness in correcting grammatical errors, which, in contrast, leads to lower ASR performance for spontaneous and colloquial speech.